# New Laboratory Constraints on Neutrinophilic Mediators

### v2

@@ II Meson decays
In the case of $\ell=e$, the logarithmic divergence is heavily suppressed by $x_{e{\sf M}}\equiv m_{e}^{2}/m_{\pi,\,K}^{2}$, hence the $C_{2}$ term becomes more important in Eq. (2) for $m_{\phi}\gtrsim$ eV. In other words, the IR divergence does not dominate the decay rate of ${\sf M}^{\pm}\to e^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$. The limits on ${\sf M}^{\pm}\to e^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$ from the ${\sf M}^{\pm}\to e^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}$ data (red lines) are flat even in the limit of $m_{\phi}\to 0$, showing little differences between the tree-level and tree+loop-level limits, as expected. When $m_{\phi}\ll m_{\sf M}$, the pion and kaon decay limits on $g_{\nu}$ in the electron channel are respectively $0.013$ and $4.3\times 10^{-3}$.
@@
By contrast, in the muon case $\ell=\mu$, as $m_{\mu}$ is comparable to $m_{\pi,K}$, the divergent behavior is noticeable in the $m_{\phi}\to 0$ limit, which is clear from the dashed blue lines in Fig. 2 (the label “tree”). The corresponding IR-free limits with the loop corrections are presented by the solid blue lines, labeled as “tree+loop”. Notably, while the tree-level decay ${\sf M}^{\pm}\to\ell^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$ is kinematically forbidden when $m_{\phi}\geq m_{\sf M}-m_{\ell}$, the 1-loop contribution to the decay ${\sf M}^{\pm}\to\ell^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}$ still exists. Consequently, the solid blue lines in Fig. 2 can extend to large $m_{\phi}$, even beyond the parent particle mass $m_{\sf M}$, whereas the dashed blue lines quickly vanish as $m_{\phi}$ gets closer to $m_{\sf M}-m_{\mu}$. With the loop contributions included, $\pi^{\pm}\to\mu^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$ gives $g_{\nu}<0.50$ in the massless $\phi$ limit. For heavy $\phi$, this decay constrains $m_{\phi}$ up to $\sim 2.5$ GeV for $g_{\nu}<1$. The IR limit for the kaon decay is relatively weaker, i.e., $g_{\nu}<1.3$ is allowed, while for the mass range of $69\,{\rm MeV}<m_{\phi}<1.65\,{\rm GeV}$ the coupling $g_{\nu}<1$. The dip feature of the solid blue line at around 200 MeV in the right panel of Fig. 2 is due to the substantial cancellation of the tree and 1-loop contributions.
These limits can be further improved by a dedicated shape analysis of the final decay products. The PIENU experiment has provided limits on the branching ratio (BR) of $\pi^{\pm}\to e^{\pm}/\mu^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+X$ as a function of the invisible $X$ mass PIENU:2021clt . We reinterpret them with our updated partial width calculations in both electron and muon channels, which are shown respectively by the orange and purple lines in the left panel of Fig. 2. In the electron channel, we find that the limit of $g_{\nu}$ is improved by $\sim 12\%$ for small $m_{\phi}$. In contrast, the loop-included result (solid) differs from the tree-level one (dashed) in the muon channel. Qualitatively, when $m_{\phi}$ approaches the kinematic threshold, $m_{\pi}-m_{\mu}\simeq 34$ MeV, the tree-level contribution is highly suppressed by phase space and the loop contribution becomes the dominant BSM effect. Therefore, the “tree+loop” limit gets much stronger at $m_{\phi}\gtrsim 10$ MeV. For $m_{\phi}\to 0$, the PIENU limit on $g_{\nu_{\mu}}$ is 0.029, while at $m_{\phi}\sim 10$ MeV, it is improved to $7.2\times 10^{-3}$.
@@
Similarly, the NA62 experiment has reported the shape-analysis-based limits on ${\rm BR}(K^{\pm}\to\mu^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+X)$ with $X$ being a scalar NA62:2021bji . They are of the order of ${\cal O}(10^{-6})$ for $37010\;{\rm MeV}<m_{X}<$ 370 MeV, roughly three orders of magnitude stronger than the limits from the partial widths above. The resultant “tree” and “tree+loop” limits are shown respectively by the dashed and solid purple lines in the right panel of Fig. 2. Again, with the 1-loop contribution included, the NA62 limits get much stronger, especially when $m_{\phi}$ is close to the threshold; at $m_{\phi}=370$ MeV, the limit can reach down to $2.1\times 10^{-3}$.
@@
Speaking of the existing limits, all the gray-shaded regions in both panels of Fig. 2 show the exclusions by current terrestrial, astrophysical, and cosmological data Berryman:2022hds , i.e., those from the cosmic microwave background (CMB) Archidiacono:2013dua ; Escudero:2019gvw ; Camarena:2024zck , big bang nucleosynthesis (BBN) Blinov:2019gcj , SN1987A Kolb:1987qy ; Shalgar:2019rqe (see also Refs. Chang:2022aas ; Fiorillo:2022cdq ; Fiorillo:2023ytr ; Fiorillo:2023cas ; Telalovic:2024cot ), IceCube High Energy Starting Events Bustamante:2020mep , and the high-energy neutrinos detected by IceCube from the neutrino point sources (NPSs) NGC 1068 and TXS 0506+056 Kelly:2018tyg ; Doring:2023vmk . Other existing limits, e.g., those from double-beta decays Deppisch:2020sqh (see also Ref. Brune:2018sab ) and stellar cooling Escudero:2019gzq , are relatively weaker for the parameter space of our interest and hence not shown in Fig. 2. More details of the limits are given in Appendix A. Moreover, the coupling in Eq. (1) induces 1-loop couplings of $\phi$ to the quarks and charged leptons Chikashige:1980ui , which would give additional limits from neutrino-electron and neutrino-nucleus scattering Laha:2013xua , e.g., those from Borexino Borexino:2008gab and COHERENT COHERENT:2017ipa . However, they are highly suppressed by the loop factor and the heavy $W$ and $Z$ particles in the loop, and are therefore not shown here.
@@ V Discussions and conclusions
In this work, the effects of the scalar $\phi$ are either via the extra decay channels of the charged mesons, the tau lepton and $Z$ boson, or via virtual correction to the wavefunction renormalization of the neutrino field. In both cases, the effects of $\phi$ can not be simply integrated out in the large mass limit, unlike in the $\phi$-induced neutrino self-scattering processes such as $\nu\nu\to\nu\nu$. Nevertheless, the limits on the $\phi$ couplings to neutrinos derived here using the high-precision SM decay processes are complementary to the direct limits on the nonstandard interactions (NSIs) of neutrinos, e.g. those from SN1987A Shalgar:2019rqe . See Appendix A for more details.
@@ Acknowledgements.
We thank Kaladi Babu, Bhaskar Dutta, Sudip Jana, Kevin Kelly, Lorenzo Ricci, Oleksandr Tomalak, Xun-Jie Xu and Yue Zhang for useful discussions and comments on the draft. BD is supported by the U.S. Department of Energy grant No. DE-SC 0017987. The work of DK is supported by the DOE Grant No. DE-SC0010813. DS is supported by NSF Grant No. PHY-2210361 and by the Maryland Center for Fundamental Physics. KS is supported by the U.S. Department of Energy grant DE-SC0009956. YZ is supported by the National Natural Science Foundation of China under grant No. 12175039, the 2021 Jiangsu Shuangchuang (Mass Innovation and Entrepreneurship) Talent Program No. JSSCBS20210144, and the “Fundamental Research Funds for the Central Universities”. BD, DK, and KS acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP* 2024) and the Institute for Underground Science at SURF for hospitality and for providing a stimulating environment, where this work was finalized.
@@ Appendix A Details of the limits
where $\tau_{\pi^{\pm}}$ and ${\rm BR}(\pi\to e\nu)$ are the expererimental values of the charged pion lifetime and the BR for the channel $\pi^{\pm}\to e^{\pm}\nu$ ParticleDataGroup:2024cfk . $f_{\pi^{\pm}}$ is the charged pion decay constant, independently determined from the lattice calculations fpi . The factor of 2 for $f_{\pi^{\pm}}$ in Eq. (4) is for the square dependence of $\Delta\Gamma(\pi\to e\nu)$ on the decay constant, cf. Eq. (2). It should be noted that in Eq. (4) the uncertainty $\Delta\Gamma(\pi\to e\nu)$ is dominated by the decay constant, whose precision is at the order of ${\cal O}(10^{-2})$ (cf. Table 1). Other more precise measurements, such as the charged pion and kaon masses, and the Fermi constant, contribute much less to the uncertainty $\Delta\Gamma(\pi\to e\nu)$, and are therefore neglected in our calculations. To apply the limits at the 90% C.L., we multiply the $1\sigma$ uncertainties by a factor of 1.64.
The SN1987A limit on $m_{\phi}$ and $g_{\nu}$ in Figs. 2, 3 and 4 is from Ref. Shalgar:2019rqe . This is from the effect of neutrino NSIs on the core of SN1987A, assuming the coupling $g_{\nu}$ is flavor diagonal and universal. This limit is more stringent than that from the scattering of the supernova neutrinos with the cosmic neutrino background en route to Earth for $m_{\phi}\gtrsim$ keV Kolb:1987qy ; Shalgar:2019rqe . For sufficiently large NSIs, neutrinos will form a fluid in the supernova core, and the outflow of neutrino fluid may behave like a burst or a steady-state wind, which occurs is however uncertain Chang:2022aas . The neutrinos might also stream as a fireball, and the effects of neutrino NSIs on supernova core might be small for some regions of the parameter space Fiorillo:2023ytr ; Fiorillo:2023cas .
@@ Appendix B Meson decay ${\sf M}\to\ell+\nu+\phi$
leads to the IR term proportional to $x_{\ell{\sf M}}(1-x_{\ell{\sf M}})^{2}\log x_{\phi{\sf M}}$, which cancels exactly with the first term in Eq. (2) from the 3-body decay ${\sf M}\to\ell+\nu+\phi$, as expected on general grounds from the KLN theorem. We emphasize that the IR and UV divergences are independent of each other, and cancellation of the UV divergence by adding counterterms does not affect the IR physics.
@@ Appendix C Tau decay $\tau\to\pi+\nu+\phi$
where the factor of $1/2$ is for averaging over the spins of tau in the initial state. Then the partial width reads

### v3

@@ II Meson decays
In the case of $\ell=e$, the logarithmic divergence is heavily suppressed by $x_{e{\sf M}}\equiv m_{e}^{2}/m_{\pi,\,K}^{2}$, hence the $C_{2}$ term becomes more important in Eq. (2) for $m_{\phi}\gtrsim$ eV. In other words, the IR divergence does not dominate the decay rate of ${\sf M}^{\pm}\to e^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$. The limits on ${\sf M}^{\pm}\to e^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$ from the ${\sf M}^{\pm}\to e^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}$ data (red lines) are flat even in the limit of $m_{\phi}\to 0$, showing little differences between the tree-level and tree+loop-level limits, as expected. When $m_{\phi}\ll m_{\sf M}$, the pion and kaon decay limits on $g_{\nu}$ in the electron channel are respectively 0.0130.0130.013 and $4.3\times 10^{-3}$.
@@
In contrast, in the muon case $\ell=\mu$, as $m_{\mu}$ is comparable to $m_{\pi,K}$, the divergent behavior is noticeable in the $m_{\phi}\to 0$ limit, which is clear from the dashed blue lines in Fig. 2 (the label “tree”). The corresponding IR-free limits with the loop corrections are presented by the solid blue lines, labeled as “tree+loop”. Notably, while the tree-level decay ${\sf M}^{\pm}\to\ell^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$ is kinematically forbidden when $m_{\phi}\geq m_{\sf M}-m_{\ell}$, the 1-loop contribution to the decay ${\sf M}^{\pm}\to\ell^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}$ still exists. Consequently, the solid blue lines in Fig. 2 can extend to large $m_{\phi}$, even beyond the parent particle mass $m_{\sf M}$, whereas the dashed blue lines quickly vanish as $m_{\phi}$ gets closer to $m_{\sf M}-m_{\mu}$. With the loop contributions included, $\pi^{\pm}\to\mu^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+\phi$ gives $g_{\nu}<0.50$ in the massless $\phi$ limit. For heavy $\phi$, this decay constrains $m_{\phi}$ up to 2.5$\sim$ 2.5∼ 2.5 GeV for $g_{\nu}<1$. The limit gets weaker in the large $m_{\phi}$ limit, as expected by the decoupling theorem. The IR limit for the kaon decay is relatively weaker, i.e., $g_{\nu}<1.3$ is allowed, while for the mass range of $69\,{\rm MeV}<m_{\phi}<1.65\,{\rm GeV}$ the coupling $g_{\nu}<1$. The dip feature of the solid blue line at around 200 MeV in the right panel of Fig. 2 is due to the substantial cancellation of the tree and 1-loop contributions.
[…]
@@
Similarly, the NA62 experiment has reported the shape-analysis-based limits on ${\rm BR}(K^{\pm}\to\mu^{\pm}+\overset{\scriptscriptstyle(-)}{\nu}+X)$ with $X$ being a scalar NA62:2021bji . They are of the order of ${\cal O}(10^{-6})$ for $10\;{\rm MeV}<m_{X}<370$ MeV, roughly three orders of magnitude stronger than the limits from the partial widths above. The resultant “tree” and “tree+loop” limits are shown respectively by the dashed and solid purple lines in the right panel of Fig. 2. Again, with the 1-loop contribution included, the NA62 limits get much stronger, especially when $m_{\phi}$ is close to the threshold; at $m_{\phi}=370$ MeV, the limit can reach down to $2.1\times 10^{-3}$.
@@
Speaking of the existing limits, all the gray-shaded regions in both panels of Fig. 2 show the exclusions by current terrestrial, astrophysical, and cosmological data Berryman:2022hds , i.e., those from the cosmic microwave background (CMB) Archidiacono:2013dua ; Escudero:2019gvw ; Camarena:2024zck , big bang nucleosynthesis (BBN) Blinov:2019gcj , IceCube High Energy Starting Events Bustamante:2020mep , and the high-energy neutrinos detected by IceCube from the neutrino point sources (NPSs) NGC 1068 and TXS 0506+056 Kelly:2018tyg ; Doring:2023vmk . Other existing limits, e.g., those from double-beta decays Deppisch:2020sqh (see also Ref. Brune:2018sab ), SN1987A Kolb:1987qy ; Shalgar:2019rqe ; Chang:2022aas ; Fiorillo:2022cdq ; Fiorillo:2023ytr ; Fiorillo:2023cas ; Telalovic:2024cot and stellar cooling Escudero:2019gzq , are relatively weaker for the parameter space of our interest and hence not shown in Fig. 2. More details of the limits are given in Appendix A. Moreover, the coupling in Eq. (1) induces 1-loop couplings of $\phi$ to the quarks and charged leptons Chikashige:1980ui , which would give additional limits from neutrino-electron and neutrino-nucleus scattering Laha:2013xua , e.g., those from Borexino Borexino:2008gab and COHERENT COHERENT:2017ipa . However, they are highly suppressed by the loop factor and the heavy $W$ and $Z$ particles in the loop, and are therefore not shown here.
@@ V Discussions and conclusions
In this work, the effects of the scalar $\phi$ are either via the extra decay channels of the charged mesons, the tau lepton and $Z$ boson, or via virtual correction to the wavefunction renormalization of the neutrino field. In both cases, the limits on the $\phi$ couplings to neutrinos derived here using the high-precision SM decay processes are complementary to the direct limits on the nonstandard interactions (NSIs) of neutrinos from SN1987A Shalgar:2019rqe , e.g., using $\phi$-induced neutrino self-scattering process $\nu\nu\to\nu\nu$. See Appendix A for more details.
@@ Acknowledgements.
We thank Kaladi Babu, Bhaskar Dutta, Damiano Fiorillo, Sudip Jana, Kevin Kelly, Lorenzo Ricci, Oleksandr Tomalak, Xun-Jie Xu and Yue Zhang for useful discussions and comments on the draft. BD is supported by the U.S. Department of Energy grant No. DE-SC 0017987. The work of DK was supported in part by the DOE Grant No. DE-SC0010813. DS is supported by NSF Grant No. PHY-2210361 and by the Maryland Center for Fundamental Physics. KS is supported by the U.S. Department of Energy grant DE-SC0009956. YZ is supported by the National Natural Science Foundation of China under grant No. 12175039, the 2021 Jiangsu Shuangchuang (Mass Innovation and Entrepreneurship) Talent Program No. JSSCBS20210144, and the “Fundamental Research Funds for the Central Universities”. BD, DK, and KS acknowledge the Center for Theoretical Underground Physics and Related Areas (CETUP* 2024) and the Institute for Underground Science at SURF for hospitality and for providing a stimulating environment, where this work was finalized.
@@ Appendix A Details of the limits
where $\tau_{\pi^{\pm}}$ and ${\rm BR}(\pi\to e\nu)$ are the experimental values of the charged pion lifetime and the BR for the channel $\pi^{\pm}\to e^{\pm}\nu$ ParticleDataGroup:2024cfk . $f_{\pi^{\pm}}$ is the charged pion decay constant, independently determined from the lattice calculations fpi . The factor of 2 for $f_{\pi^{\pm}}$ in Eq. (4) is for the square dependence of $\Delta\Gamma(\pi\to e\nu)$ on the decay constant, cf. Eq. (2). It should be noted that in Eq. (4) the uncertainty $\Delta\Gamma(\pi\to e\nu)$ is dominated by the decay constant, whose precision is at the order of ${\cal O}(10^{-2})$ (cf. Table 1). Other more precise measurements, such as the charged pion and kaon masses, and the Fermi constant, contribute much less to the uncertainty $\Delta\Gamma(\pi\to e\nu)$, and are therefore neglected in our calculations. To apply the limits at the 90% C.L., we multiply the $1\sigma$ uncertainties by a factor of 1.64.
Neutrino NSIs may potentially have impacts on the supernova core. For sufficiently large NSIs, neutrinos will form a fluid in the supernova core, and the outflow of neutrino fluid streams as a fireball Chang:2022aas . Therefore, the impact of large neutrino NSIs on supernova dynamics is small Fiorillo:2023ytr ; Fiorillo:2023cas . There are also constraints from the scattering of the supernova neutrinos with the cosmic neutrino background en route to Earth for $m_{\phi}\gtrsim$ keV. However, the corresponding limits are rather weak for the parameter space of $m_{\phi}$ and $g_{\nu}$ we are interested in Kolb:1987qy ; Shalgar:2019rqe .
@@ Appendix B Meson decay ${\sf M}\to\ell+\nu+\phi$
leads to the IR term proportional to $x_{\ell{\sf M}}(1-x_{\ell{\sf M}})^{2}\log x_{\phi{\sf M}}$, which cancels exactly with the first term in Eq. (2) from the 3-body decay ${\sf M}\to\ell+\nu+\phi$, as expected on general grounds from the KLN theorem. We emphasize that the IR and UV divergences are independent of each other, and cancellation of the UV divergence by adding counterterms should not affect the IR behavior.
@@ Appendix C Tau decay $\tau\to\pi+\nu+\phi$
where the factor of 1/21/21 / 2 is for averaging over the spins of tau in the initial state. Then the partial width reads